OpenStudy (anonymous):

there are three ants on the corners of a triangle nd they start moving randomly towards one another.......wat's the probability that they dont collide

OpenStudy (anonymous):

could you define "moving randomly towards one another"?

OpenStudy (anonymous):

Each of them can move in 2 directions. So number of case is 2*2*2 = 8 They dont collide if all move in clockwise or counter clockwise. So number of that case is 2. I think the pr. here is 2/8 = 1/4

OpenStudy (anonymous):

i am going to hazard a guess the question is slightly different than posed, because as it is i think you would need lots more information (size of the ants, for one)

OpenStudy (anonymous):

@jeanlouie answered the question "There are 3 ants at 3 corners of a triangle they randomly start moving towards another corner what is the probability that they do not collide?" which is not the one posted, but probably the actual question

OpenStudy (anonymous):

sorry @satellite73 that's the entire question and answer is 25 % r say 1/4

OpenStudy (anonymous):

big difference right? randomly moving towards ANOTHER CORNER versus randomly moving towards EACH OTHER

OpenStudy (anonymous):

i m sorry i dint xplain the question properly .... really sorry.... it means corner......

OpenStudy (anonymous):

yea it is ...it is lot more complicated if it is towards each other....... can you solve it @satellite73

OpenStudy (anonymous):

@jeanlouie has it,

OpenStudy (anonymous):

@satellite73 oh your idea is really interesting

OpenStudy (anonymous):

must be a mind reader

hero (hero):

Saying "I'm lost" would be an understatement

OpenStudy (anonymous):

easy fellas

OpenStudy (anonymous):

@jeanlouie has the correct answer and correct reasoning